solve. approximate irrational roots to the nearest tenth.
The length of a rectangle is 3 cm greater than its width. The area of the rectangle is 180 cm^2. find the length and width. (units optional)

Answer by adamchapman(301)

W = width
L = length
A = area = 180
The length of the rectangle is 2 inches more than twice its width:
L = W + 3..............(i)
A = W x L..............(ii)
Subtitute equation (i) into (ii) for L:
A = W(W + 3)
A = W^2 + 3W
180 = W^2 + 3W
W^2 + 3W -180 = 0
Solve the quadratic equation to give W=-12.75cm or W=14.25cm
Obviously the first solution is impossible, so the true answer is W=14.25cm. Add 3cm to W to get L=17.25cm
I hope this helps
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk